A ball of mass 50g is hit due south with a speed of 70 ms-1 at latitude 60 ºN.
Calculate the magnitude and direction of the coriolis force on the ball.
1
Expert's answer
2020-05-14T09:28:40-0400

The magnitude and direction of the Coriolis force on the ball:


"F_c=2mvΩ\text{ sin}\phi=\\
=2mv\frac{2\pi R_e}{T_e}\text{ sin}\phi=\\
=2\cdot0.05\cdot70\cdot\frac{2\pi (6375\cdot10^3)}{24\cdot3600}\text{ sin}60^\circ=4.4\cdot10^{-5}\text{ N}"


due east.